The financial statements of Apple Inc. show the following numbers for fiscal year 2017 ($000,000): Revenue...

  1. The financial statements of Apple Inc. show the following numbers for fiscal year 2017 ($000,000):

Revenue (Net Sales)

$229,234

Net Income

$48,351

Total Assets (Year end 2017/2016)

$348,503

Shareholder Equity

$134,047

Use the Dupont Equation to show the firm’s return on equity (ROE) and the drivers of each component in the equation (Return on Sales, Return on Assets, and Leverage) in determining this number.
Answer #1

ROE = net profit margin*total asset turnover*equity multiplier

Net profit margin = net income/sales = 48351/229234=21.09%

Total asset turnover = sales/total assets = 229234/348503=0.6577

Equity multiplier = asset/equity = 348503/134047=2.5999

therefore

ROE = 21.09*0.6577*2.5999=36.06%